Can someone help on this problem?
The revenue for a sandwich shop is directly proportional to ites advertising budget. When the owner spent $2,000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner spending on advertising?
Thanks, Ashley

Found 2 solutions by stanbon, Genius:
Answer by stanbon(48545)

The revenue for a sandwich shop is directly proportional to ites advertising budget. When the owner spent $2,000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner spending on advertising?
--------
rev = k(adbud)
Need to find "k"
120,000= k(2000)
k=60
So, rev = 60(adbud)
180,000=60(adbud)
adbud= 180000/60= $3000
Cheers,
Stan H.

Answer by Genius(17)

2,000/120,000*180,000=
0.016666666666666666666666666666667*180,000=3,000
$3,000